The length of a rectangular garden is 8 ft less than 2 times its width. If the perimeter of the garden is 44 ft, find the dimensions of the garden.

Answer by Paul(988)

Let the width be x
Let the length be 2x-8
together:
2(x+2x-8)=44
6x-16=44
6x=60
x=10
2(10)-8=12
Hence, the length is 12ft and the width is 10ft.
Paul.